OpenStudy (anonymous):

It was proved that D(xr)=rxr−1 for any positive integer r. Use this result and the Quotient Rule to prove that D(xr)=rxr−1 for any negative integer r.

OpenStudy (anonymous):

Quotient rule: \[D \left( \frac{ f }{ g } \right)=\frac{ g*f' - f*g' }{ g² }\]

terenzreignz (terenzreignz):

So... let r be a positive integer... so -r is a negative integer, aye? :)

terenzreignz (terenzreignz):

On second thought, this is confusing... Let's let r be a negative integer instead...

OpenStudy (anonymous):

\[D(x^r)=rx^r⁻¹\]

OpenStudy (anonymous):

sorry the first post looks confusing

terenzreignz (terenzreignz):

Yeah... so... any idea how to go about this?

OpenStudy (anonymous):

I'm thinking how to get two functions g and f out of the original function

terenzreignz (terenzreignz):

Well, you'll find that most work in Maths involves reducing a problem into a mess of previously solved problems... So now we're to find the derivative of x^r, given that r is a negative integer, aye? So... \[\huge \frac{d}{dx}x^r \ \ \ \ \ r\in\mathbb{Z}^-\]

OpenStudy (anonymous):

so... I can use the power rule, but that doesn't really prove anything new?

OpenStudy (anonymous):

x^r= 1/x^-r ........... so now I have quotient?

OpenStudy (anonymous):

...when r is a negative integer ofcourse

terenzreignz (terenzreignz):

You seem to have a good idea how to do it :) So, now you have \[\huge \frac{d}{dx}x^r=\frac{d}{dx}\frac1{x^{-r}}\]And so you have a quotient... ripe for quotient rule :D Get to it, champ :)

terenzreignz (terenzreignz):

Not forgetting of course, that since r was negative, -r is positive, so you can use the power rule...

OpenStudy (anonymous):

\[D \left( \frac{ 1 }{ x^-r } \right)=\frac{ x⁻^r*0-1*r*x⁽⁻^r⁻¹⁾ }{ x^-²^r }\] ?

OpenStudy (anonymous):

whoops missing a minus sign on the numerator

terenzreignz (terenzreignz):

Rectify it :)

OpenStudy (anonymous):

\[\frac{ -1*-rx⁻^r⁻¹ }{ x⁻²^r }=\frac{ -1*-r*\left( \frac{ x⁻^r }{ x¹ } \right) }{ x⁻^r*x⁻^r }\] is this right?

terenzreignz (terenzreignz):

Hang on... let me just zoom it in :D \[\huge \frac{ -1*-rx⁻^r⁻¹ }{ x⁻²^r }=\frac{ -1*-r*\left( \frac{ x⁻^r }{ x¹ } \right) }{ x⁻^r*x⁻^r }\]

terenzreignz (terenzreignz):

Yes, it's right. Now simplify it.

OpenStudy (anonymous):

Ok I can basically see it will become rx^(r-1), but I'm not sure what the rules are for dividing quotients, or simplifying them in this situation

terenzreignz (terenzreignz):

Dividing quotients? What do you mean by that?

OpenStudy (anonymous):

when there is \[\left( \frac{ x⁻^r }{ x¹ } \right) / (x⁻^r*x⁻^r)\]

OpenStudy (anonymous):

what cancels out what

OpenStudy (anonymous):

can I multiply by the inverse?

terenzreignz (terenzreignz):

Ahh... Well... Given this... \[\huge \frac{\frac{a}{b}}{c}=\frac{a}{bc}\]should clear it up.

OpenStudy (anonymous):

\[=\frac{ r*x⁻^r }{ x*x⁻^r*x⁻^r }=\frac{ r }{ x*x⁻^r }=\frac{ r }{ x⁻^r⁺¹ }\] ??

OpenStudy (anonymous):

=\[\frac{ r }{ x^r⁻¹ }\]

terenzreignz (terenzreignz):

This is good. You've done it :)

terenzreignz (terenzreignz):

If you doubt yourself, do remember one thing about exponents... \[\huge a^m=\frac1{a^{-m}}\]

OpenStudy (anonymous):

wait, but how do I get to the result rx^r-1 ?

terenzreignz (terenzreignz):

Oh wait...

terenzreignz (terenzreignz):

You stop at... \[\huge =\frac{ r\times x⁻^r }{ x\times x⁻^r\times x⁻^r }=\frac{ r }{ x\times x⁻^r }=\frac{ r }{ x⁻^r⁺¹ }\]

terenzreignz (terenzreignz):

The last part has... \[\huge \frac{r}{x^{-r+1}}\]

terenzreignz (terenzreignz):

Right?

OpenStudy (anonymous):

aye

terenzreignz (terenzreignz):

Well, it's the same as... \[\huge \frac{r}{x^{1-r}}\]and you can see it now, right? :)

terenzreignz (terenzreignz):

Actually, much more better would be... \[\huge \frac{r}{x^{-r+1}}=\frac{r}{x^{-(r-1)}}\]

terenzreignz (terenzreignz):

You know where this is going? :)

OpenStudy (anonymous):

y^-z=1/y^z works in reverse too?

terenzreignz (terenzreignz):

Of course :)

OpenStudy (anonymous):

wow thanks a lot for helping me solve this one!

terenzreignz (terenzreignz):

No problem :) You practically solved it yourself, anyway :D